OpenStudy (mathlegend):

Solve Radical Equation:

OpenStudy (mathlegend):

\[\sqrt{1-x ^{2}}+2=3\]

OpenStudy (nincompoop):

you're kidding right?

OpenStudy (mathlegend):

@nincompoop actually I am not... it has been awhile since I have looked at math this summer.

OpenStudy (nincompoop):

how do you want to do this?

OpenStudy (nincompoop):

get rid of the radical first?

OpenStudy (mathlegend):

I was going to subtract two from 3.

OpenStudy (anonymous):

First bring the 2 on to the other side: sqrt(1-x^2) = 1 Now square both sides: 1-x^2 = 1 Can you solve it from here? Btw at the end you will get 2 values for 'x'. Plug them both in to the original equation to see which one ACTUALLY works. @MathLegend

OpenStudy (anonymous):

Yes, that will be your first step.

OpenStudy (nincompoop):

yes

OpenStudy (nincompoop):

correct then?

OpenStudy (mathlegend):

Then I can square the 1-x^2 and the 1 on the opposite side of the equal sign?

OpenStudy (nincompoop):

okay good and then?

OpenStudy (mathlegend):

Yeah but I'm left with... 1-x^2 = 1 ???

OpenStudy (nincompoop):

okay and then? you want x right? keep isolating it

OpenStudy (mathlegend):

add 1 to both sides

OpenStudy (nincompoop):

add?

OpenStudy (mathlegend):

subtract

OpenStudy (nincompoop):

good and then?

OpenStudy (mathlegend):

-x^2 = 0

OpenStudy (nincompoop):

then?

OpenStudy (mathlegend):

do I square root this?

OpenStudy (anonymous):

After you subtract, what is left?

OpenStudy (mathlegend):

-x^2=0

OpenStudy (nincompoop):

you want to remove -1 first remember that -x^ means -1x^2

OpenStudy (mathlegend):

divide by -1?

OpenStudy (nincompoop):

okay do it

OpenStudy (mathlegend):

x^2 = 0

OpenStudy (nincompoop):

then undo the ^2 by?

OpenStudy (anonymous):

Yes, because you cannot take a square root of a negative number.

OpenStudy (mathlegend):

square root it.. but the square root of 0 is 0 right?

OpenStudy (anonymous):

You got it!!!

OpenStudy (mertsj):

1. Subtract 2 from both sides. 2. Square both sides. 3. Subtract 1 from both sides. 4. Multiply both sides by -1 5. Take the square root of both sides.

OpenStudy (nincompoop):

see? you're not that rusty when you rest you rust

OpenStudy (mathlegend):

:) I hope I can finish this packet...

OpenStudy (anonymous):

You'll do fine. You have been doing great so far. I think your biggest obstacle is not being confident in your knowledge. You do seem to know what you are doing. :-)

OpenStudy (nincompoop):

PRACTICE ON THIS TOO OR REHASH YOUR KNOWLEDGE FROM IT http://www.scribd.com/doc/72151240/Michael-Spivak-Calculus